Given that sin@ = 3/5 and cos B = 12/13 ,prove that one possible value of (@+B) is 33/65 ,and find all other possible values?
Found 3 solutions by Alan3354, ikleyn, greenestamps:
Answer by Alan3354(69443)

Given that sin@ = 3/5 and cos B = 12/13 ,prove that one possible value of (@+B) is 33/65 ,and find all other possible values?
===========
I suspect your entry is not correct.

Answer by ikleyn(52835)

.

I am sure that your entry is wrong.



Answer by greenestamps(13203)



Your post is almost certainly not what you intended....

"(@+B)" is an angle; "33/65" looks like the sine of an angle or the cosine of an angle. It is highly unlikely that "33/65" would be the measure of an angle.

I could guess what the problem is supposed to say, and answer that. But I shouldn't have to do that.

It is your responsibility, if you are serious about getting help with your problems, to take care to post them correctly.

Try again....